# Training-Conditional Coverage Bounds under Covariate Shift

### v2

@@ 3.1 Split conformal
where $m^{\prime}=m\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-m^{1/2+\epsilon}$.
@@ 3.2 Full conformal and jackknife+
4 Fully conditional conformal prediction
In this section, we study conformal prediction conditional on both training dataset and test bin.
@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
where $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-n^{1/2+\epsilon}$.
@@
where $n_{B}=\sum_{i\in\mathcal{I}}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\}$. Now let $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-n^{1/2+\epsilon}$ and observe that
@@
$=\operatorname{\mathbb{E}}\left(\left(1-4e^{-n_{B}\delta^{2}/(2B^{2})}\right)\boldsymbol{1}\{n_{B}\geq n^{\prime}\}\right)$
$\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\operatorname{\mathbb{P}}\left(n_{B}\geq n^{\prime}\right)\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\left(1-2e^{-2n^{2\epsilon}}\right)$
$\geq 1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}-2e^{-2n^{2\epsilon}}.$
@@
$|{\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)|$ $=\left|\frac{W^{<}(x)}{W^{<}}-\frac{W^{<}(x)+W^{>}(x)}{W^{<}+W^{>}}\right|$
@@
$\underset{x\in\mathbb{R}}{\sup}\ |{\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)|\leq 2\ \frac{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}\overset{a.s.}{\to}2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right).$
@@
$\leq\delta+2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)$
with probability at least Durrett [2019][Theorem 2.2.12],
$1-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)}.$
Thus,
@@
$1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}-2e^{-2n^{2\epsilon}}-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)}.$
Remark 2
Observe that ${\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)\overset{a.s.}{\to}{F}_{Q|\mathcal{X}_{B}}(x)-{F}_{Q}(x)$. Also note that ${\hat{F}}_{Q,m|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,m}(x)\overset{a.s.}{\to}0$ if $s(Z_{\text{test}})$ and $\frac{dQ}{dP}(Z_{\text{test}})$ are independent.

### v3

@@ 3.1 Split conformal
where $m^{\prime}=m\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{calib}})\leq B\right)-m^{1/2+\epsilon}$.
@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
where $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{calib}})\leq B\right)-n^{1/2+\epsilon}$.
@@
where $n_{B}=\sum_{i\in\mathcal{I}}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})\leq B\right\}$. Now let $n^{\prime}=n\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{calib}})\leq B\right)-n^{1/2+\epsilon}$ and observe that
@@
$\geq\operatorname{\mathbb{E}}\left(\left(1-4e^{-n_{B}\delta^{2}/(2B^{2})}\right)\boldsymbol{1}\{n_{B}\geq n^{\prime}\}\right)$
$\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\operatorname{\mathbb{P}}\left(n_{B}\geq n^{\prime}\right)\overset{(*)}{\geq}\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\left(1-2e^{-2n^{2\epsilon}}\right)$
$\geq 1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}-2e^{-2n^{2\epsilon}},$
where ()(*)( ∗ ) holds according to Hoeffding’s inequality.
@@
$|{\hat{F}}_{Q,n|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,n}(x)|$ $=\left|\frac{W^{<}(x)}{W^{<}}-\frac{W^{<}(x)+W^{>}(x)}{W^{<}+W^{>}}\right|$
@@
$\underset{x\in\mathbb{R}}{\sup}\ |{\hat{F}}_{Q,n|\mathcal{X}_{B}}(x)-{\hat{F}}_{Q,n}(x)|\leq 2\ \frac{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{i})>B\right\}}{\sum_{i\in\mathcal{I}}\frac{dQ}{dP}(Z_{i})}\overset{a.s.}{\to}2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right).$
@@
$\leq\delta+2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right),$
with probability at least
$1-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)},$
where in the last step we have used (the proof of) Durrett [2019][Theorem 2.2.11] with $\epsilon=\delta/2$ and $b_{n}=n$ twice; once for the first term with $X_{n,k}=\frac{dQ}{dP}(Z_{k})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{k})>B\right\}$, $k\in\mathcal{I}$, and the other time for the third term with $X_{n,k}=dQ/dP(Z_{k})$, $k\in\mathcal{I}$, and the requirements of the theorem are satisfied by $\operatorname{\mathbb{E}}dQ/dP(Z_{\text{calib}})=1<\infty$ according to Durrett [2019][Theorem 2.2.14]; also recall that $\mathcal{I}\subset[n]$ for the split conformal method corresponds to the calibration dataset. Putting everything together, we have
@@
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)}.$